OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x − 1)? A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 greater than or equal to x x less than or greater to 3

OpenStudy (mathteacher1729):

Here is a step-by-step breakdown. Start at the top and read down the line, making sure you understand each step. \[ \begin{align*} 5(x + 2) &\geq 7x + 2(x − 1) \quad\text{ distribute the 5 and the 2}\\ 5x +10 &\geq 7x + 2x -2 \quad\text{ combine $7x$ and $2x$}\\ 5x+10&\geq 9x -2 \quad\text{ move all constants to the left of the $\geq$ and all variables to the right}\\ 10+2&\geq 9x-5x \quad\text{ combine variables and constants.}\\ 12 &\geq 4x\quad\text{divide both sides by 4} \\ 3 &\geq x \end{align*} \]

OpenStudy (anonymous):

OMG THANK YOU< YOU SAVED ME!